This doesn't make sense...Someone help me please!The equation is:
r^3=2r^2+15r

Found 2 solutions by lynnlo, tinbar:
Answer by lynnlo(4176)

r^3=2r^2+15r
r^3-2r^2-15r=0
(r-5)r(r+3)=0
r-5=0,r=0,r+3=0
r=5,r=0,r+3=0

Answer by tinbar(133)

r^3=2r^2+15r
r^3 - 2r^2 - 15r = 0
r(r^2 - 2r - 15) = 0
Therefore r = 0, is one solution and whenever (r^2 - 2r - 15) = 0, the equation is satisfied.
r^2 - 2r - 15 = 0
(r-5)(r+3) = 0
Therefore, r = 5 and r = -3 are also solutions.
Finally, r = 0, r = -3, and r = 5 are the solutions to the original equation.